Two 60 cm2 parallel plates are connected to a battery that It provides a voltage of 12 volts.
The plates are separated by a 1.0 mm thick air layer. If the space between the plates is filled with oil and disconnect the battery
(isolated capacitor): K for oil is 4.
a) How much is capacitance?
b) How much is the load? Explain
c) How much is the voltage across the trainer?
d) How much is the electric power stored?
Answer #1

a)

A = area of plate = 60 cm2 = 60 x 10-4 m2

d = separation between the plates = 1 mm = 1 x 10-3 m

Capacitance when there is air between the plates is given as

C = A/d

C = (8.85 x 10-12) (60 x 10-4)/(1 x 10-3)

C = 5.31 x 10-11 F

Capacitance when there is oil between the plates is given as

C' = kA/d

C' = (4) (8.85 x 10-12) (60 x 10-4)/(1 x 10-3)

C' = 21.24 x 10-11 F